The manager of a local monopoly estimates that the elasticity of demand for its products is constant and equal to -3. The firm's marginal cost is constant at $35 per unit.

a. express the firm's revenue as a function of its price

MR= ???x P

b. Determine the profit-maximizing price
Answer #1

Ans) 1) Total revenue = Price × quantity

Marginal revenue = P [1+(1/Ep)] ; where Ep is elasticity.

MR= P [1+(1/-3)] = 2/3 P (put elasticity with sign)

2) Monopoly maximises profit where MR = MC

2/3 P = 35

P = (35×3)/2

P = $52.5